### v1

# MnLargeSymbols’164 MnLargeSymbols’171

Abstract.
Secure collaborative analytics (SCA) enable the processing of analytical SQL queries across multiple owners’ data, even when direct data sharing is not feasible. Although essential for strong privacy, the large overhead from data-oblivious primitives in traditional SCA has hindered its practical adoption. Recent SCA variants that permit controlled leakages under differential privacy (DP) show a better balance between privacy and efficiency. However, they still face significant challenges, such as potentially unbounded privacy loss, suboptimal query planning, and lossy processing.
To address these challenges, we introduce SPECIAL, the first SCA system that simultaneously ensures bounded privacy loss, advanced query planning, and lossless processing. SPECIAL employs a novel synopsis-assisted secure processing model, where a one-time privacy cost is spent to acquire private synopses (table statistics) from owner data. These synopses then allow SPECIAL to estimate (compaction) sizes for secure operations (e.g., filter, join) and index encrypted data without extra privacy loss. Crucially, these estimates and indexes can be prepared before runtime, thereby facilitating efficient query planning and accurate cost estimations. Moreover, by using one-sided noise mechanisms and private upper bound techniques, SPECIAL ensures strict lossless processing for complex queries (e.g., multi-join). Through a comprehensive benchmark, we show that SPECIAL significantly outperforms cutting-edge SCAs, with up to 80×80$\times$80 × faster query times and over 900×900$\times$900 × smaller memory for complex queries. Moreover, it also achieves up to an 89×89$\times$89 × reduction in privacy loss under continual processing.
1. Introduction
Organizations, such as hospitals, frequently hold sensitive data in separate silos to comply with privacy laws, despite the valuable insights that could be gained from sharing this information. Recent advancement of Secure Collaborative Analytics (SCA) (Eskandarian and Zaharia, 2017; Qin et al., 2022; Roy Chowdhury et al., 2020; Wang et al., 2021, 2022; Bater et al., 2017, 2018; Liagouris et al., 2023; Poddar et al., 2021; Bater et al., 2020; He et al., 2015) provides an exciting solution to tackle this dilemma. These systems leverage advanced multi-party secure computation (MPC) (Yao, 1986) primitives to empower multiple data owners, who previously could not directly share data, to collaboratively process analytical queries over their combined data while ensuring the privacy of each individual’s data.
While MPC can effectively conceal data values (Yao, 1986), its security guarantees do not immediately extend to the protection of execution transcripts. Consequently, data-dependent processing patterns such as memory traces and read/write volumes can still reveal critical information, risking privacy breaches (Cash et al., 2015; Kellaris et al., 2016; Blackstone et al., 2019; Oya and Kerschbaum, 2021; Grubbs et al., 2018; Shang et al., 2021; Zhang et al., 2016) even when the core data remains encrypted. To ensure strong privacy, modern SCA systems also utilize data-oblivious primitives that exhaustively pad query processing complexities to a worst-case, data-independent upper bound (Bater et al., 2017; Liagouris et al., 2023; Poddar et al., 2021). However, such stringent protections can largely reduce system efficiency and hinder the generalization of conventional optimization techniques to SCA, which are typically data-dependent (Liagouris et al., 2023). To address this, recent efforts (Qin et al., 2022; Wang et al., 2021, 2022; Bater et al., 2018) have introduced differentially private SCA (DPSCA). This approach allows controlled information leakage under differential privacy (Dwork et al., 2014) to mitigate constant worst-case overhead. For instance, systems under this model can dynamically compact intermediate query sizes to a noisy estimate close to actual sizes, avoiding exhaustive padding. As such, queries under DPSCA experience largely boosted efficiencies (e.g., up to $10^{5}\times$ faster (Wang et al., 2022)) compared to their “no leakage” counterparts. Despite these substantial performance gains, existing DPSCAs still face critical limitations that impede their practical uses, as elaborated below:
$\bullet$ L-1. Unbounded privacy loss. Most DPSCA systems utilize a per-operator privacy expenditure model (Bater et al., 2018; Wang et al., 2021; Qin et al., 2022; Chu et al., 2021; Chan et al., 2022; Wang et al., 2022), meaning each query operator (e.g., join, filter) independently consumes a portion of the privacy budget. This approach can lead to either unbounded privacy loss or the forced cessation of query responses upon budget exhaustion. To mitigate this, some studies (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., [n.d.]) propose private, locality-sensitive grouping, incurring a one-time privacy cost to pre-group data based on specific attributes. Subsequent queries on those attributes can be directly applied to a smaller subset and need no additional privacy budget. However, this method only supports simple queries (e.g., point and range); complex queries like joins still suffer from unbounded privacy loss.
$\bullet$ L-2. Suboptimal execution plan. Conventional query planners can pre-estimate sizes for equivalent plans of a given query and select the most efficient plan with minimized intermediate sizes before execution (Silberschatz et al., 2011; Blasgen et al., 1981). In contrast, SCA systems lack this capability, and even DPSCA designs (Qin et al., 2022; Wang et al., 2022; Zhang et al., 2023; Wagh et al., 2021; Bater et al., 2018, 2020) can only reactively determine plan sizes during runtime. This inherent limitation often forces existing systems to settle for less efficient query plans, such as suboptimal join orders, which lead to significantly inflated intermediate sizes (§ 7.2) and substantially hinder performance.
$\bullet$ L-3. Lossy processing. Noise from randomized mechanisms in DPSCA also introduces a unique accuracy issue (e.g., losing tuples after compaction), and unfortunately, no existing design can provide deterministic accuracy guarantees for complex queries (Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Zhang et al., 2023; Groce et al., 2019). Furthermore, stronger privacy settings can further increase noise variance, which amplifies errors, significantly impacting the utility of SCA systems.
1.1. Overview of SPECIAL.
In this work, we introduce SPECIAL, an innovative SCA system designed to address all these challenges simultaneously through a new paradigm called synopsis-assisted secure processing. At its core, SPECIAL incurs a one-time privacy cost to gather DP synopses (statistics of base tables) from owners’ data. These synopses are then used to compact query intermediate results, index encrypted data, and enhance SCA query planning. Notably, SPECIAL is the first system to provide all of the following benefits: (1) Bounded privacy—it manages complex queries, such as multi joins, within strict privacy limits; (2) Advanced query planning—it builds an advanced SCA planner that can exploit plan sizes before runtime; and (3) Lossless processing—it ensures exact results with no data omissions. An overview of SPECIAL’s architecture is shown in Figure 1.
Figure 1. Overview of SPECIAL workflow.
SPECIAL operates under a standard server-aided MPC model (Kamara et al., 2011) with three key participants: data owners, at least two SPECIAL servers, and a vetted analyst. The process starts with data owners securely outsourcing their data (e.g. using secret sharing (Kamara et al., 2011; Wang et al., 2021; Zhang et al., 2023; Wang et al., 2022)), and privately release a set of DP synopses (table statistics) to the servers. Once the data and synopses are ready, analysts can then submit queries on the outsourced data. To process queries, SPECIAL uses synopses to privately estimate tight result compaction bounds or build private indexes for faster access, thus avoiding exhaustive padding. Crucially, this information is exploit before runtime, so that SPECIAL can pre-identify an optimal secure execution plan that minimizes intermediate sizes and overall costs. Finally, the optimized plan is executed via MPC across the secret-shared data, and the results are securely returned to the analyst.
In the SPECIAL prototype, we adopt the same privacy model used in existing DPSCA designs, assuming a computationally bounded, “honest-but-curious” adversary (Goldreich, 2004) who can compromise up to $n-1$ out of $n$ owners and $m-1$ out of $m$ servers. SPECIAL ensures that such an adversary can only learn a limited leakage profile $\mathsf{Lkg}$ about honest parties’ data, which strictly adheres to DP constraints. Note that, SPECIAL’s core components, including synopses management and query planning do not leverage raw data at all. Thus, any of the servers, alone or jointly, can manage these tasks. Also, once analysts are authorized by data owners, they are considered trusted, and additional randomization of query results is not required. It is worth mentioning that SPECIAL’s corruption model (e.g. all except one) is entirely determined by the underlying MPC, and SPECIAL’s design does not modify them at all. Hence, one may also opt to a weaker corruption model, such as a supermajority of owners and servers must remain uncorrupted (Liagouris et al., 2023; Tan et al., 2021), to enhance efficiency.
1.2. Unique challenges and key contributions
Leveraging DP synopses in SCA holds significant promise for achieving our desired objectives. However, this also introduces unique challenges. Below, we highlight the key challenges and summarize our non-trivial contributions to address them:
$\bullet$ C-1. How to select proper synopses? Even a single relation can have numerous options for generating synopses, such as using different attributes or their combinations. However, improper selection can lead to large errors (e.g. using too many synopses or leverage high-dimensional attributes (Zhang et al., 2021)), or reduced functionalities (e.g., using only simple attributes (Zhang et al., 2023)). Hence, a key challenge is selecting a limited set of DP synopses to optimize the privacy budget for complex query processing. Our approach is informed by two observations: (i) secure joins are resource-intensive and need prioritized acceleration, and (ii) synopses for common filtering predicates are vital as they allow pre-built indexes on base relations for fast access. Consequently, we propose a focused strategy (§ 4) that targets low-dimensional (1D and 2D) attributes frequently involved in joins and filters within a representative workload.
$\bullet$ C-2. How to enforce lossless processing? Private synopses do not immediately implies lossless guarantees. Thus, a second challenge is designing practical approaches to achieve lossless results without violating privacy goals. To address this, we employ one-sided DP noise (either strictly positive or negative, § 4) for generating synopses and design novel primitives (§ 5) based on them to pessimistically estimate filter cardinalities and intervals of index structures. This enable us to efficiently discard unnecessary records during processing without breaking privacy and lossless guarantees. To ensure lossless processing of complex joins, we extend upon cutting-edge join upper bound techniques (Hertzschuch et al., 2021) to privately estimate join compaction sizes using available synopses. To our knowledge, this is the first study to support private join upper bound estimation.
$\bullet$ C-3. How DP synopses can empower efficient query processing? The use of DP synopses in SCA is largely underexplored, leaving a knowledge gap regarding their potential to enhance query efficiency. To navigate this potential, we explore various use of synopses in accelerating secure processing including private indexes SPEidx (§ 5.2), and compacted oblivious operations SPEop (§ 5.3). We then develop a Selinger-style query planner (§ 6) that supports both standard secure operators and SPEop. Employing a customized cost model and tailored heuristics, this planner efficiently generates optimal execution plans for queries using a mix of standard and SPEop operators. This in essence, also address the limitations of sub-optimal execution plans inherent in standard DPSCA.
$\bullet$ C-4. How to systematically evaluate SPECIAL? A major challenge in SCA design evaluation is the absence of open-source benchmarks. We address this by initiating an open-source evaluation platform accessible to the public, utilizing public Financial data (Rel, [n.d.]) (1.1M rows and 59 columns), and designing eight test queries ranging from simple linear to complex 5-way joins. We evaluated our prototype SPECIAL alongside the state-of-the-art DPSCA system, Shrinkwrap (Bater et al., 2018), and conventional SCA, SMCQL (Bater et al., 2017). Results indicate that SPECIAL outperforms Shrinkwrap, reducing query latency by up to 80.3×80.3$\times$80.3 ×, and SMCQL, with at least a 114×114$\times$114 × reduction in query latency. Additionally, SPECIAL improves memory efficiency in complex join processing by more than 900×900$\times$900 × compared to both systems. Moreover, scaling experiments show that SPECIAL can effectively scale linear and binary joins up to 8×8$\times$8 × (8.8M rows) and 5-way joins up to 4×4$\times$4 × (4.4M rows) workloads. All benchmarks, including our prototype implementation, are open-sourced and available at (Lovingmage, 2024).
The remainder of the paper is structured as follows: We begin with essential background in § 2 and a formal definition of our privacy model in § 3 before delving into design details. § 4 discusses the SPECIAL catalog design, including synopses selection and generation. § 5 details novel synopses-assisted secure database operators, and § 6 describes our planner design. Benchmark evaluations are presented in § 7, followed related works and conclusions.
2. Background
General notations. We consider the logical database $\mathcal{D}$ to contain multiple private relations $\{D_{1},D_{2},...\}$, where each relation $D_{i}$ is belongs owned by a specific party $P_{i}$, where $D\in\mathcal{D}$ is owned by a specific owner and has a set of attributes as $attr(D)$. The domain for attribute $A\in attr(D)$ is given by $dom(A)$, and the combined domain for a collection of attributes $\mathbf{A}=\{A_{1},A_{2},...\}\subseteq attr(D)$ is denoted as $dom(\mathbf{A})=\prod_{A\in\mathbf{A}}dom(A)$. For a tuple $t\in D$, and $\mathbf{A}\subseteq attr(D)$, we denote $t.\mathbf{A}$ as the attribute value of $\mathbf{A}$ in $t$
Frequency (count). Given $D$, and $\mathbf{A}\subseteq attr(D)$, and $\mathbf{v}\in dom(\mathbf{A})$, the frequency (count) of $\mathbf{v}$ in $D$ is the total number of tuples $t\in D$ with $t.\mathbf{A}=\mathbf{v}$. In addition, the max frequency moments (MF) of $\mathbf{A}$ is defined as ${\mathtt{mf}}(\mathbf{A},D)=\max_{\mathbf{v}\in dom(\mathbf{A})}|\{t\in D\mid t.\mathbf{A}=\mathbf{v}\}|$.
Histograms. Given $D$, and $\mathbf{A}\subseteq attr(D)$, the (equal-width) histogram $\mathbf{h}(\mathbf{A},D)=(c_{1},c_{2},...,c_{m})$ is a list of counts for the attribute values in $\mathbf{A}$. Specifically, $\mathbf{h}$ partitions [^1] We use a condensed notation that omits arguments when $\mathbf{A}$ and $D$ are obvious. $dom(\mathbf{A})$ into $m$ “equal-sized” domain intervals $(B_{1},...,B_{m})$, and each count $c_{i}\in\mathbf{h}$ corresponds to the total number of $t\in D$ such that $t.\mathbf{A}$ fall within $B_{i}$.
Multi-party secure computation (MPC). MPC (Yao, 1986; Goldreich, 2009; Micali et al., 1987; Ben-Or et al., 2019) is a cryptographic techniques to allow a group of participants $P_{1},P_{2},...$ to jointly compute a function $f(x_{1},x_{2},...)$ over their respective private input $x_{i}$. MPC ensures that no unauthorized information is revealed to any involved party, excepted the desired output of function $f$, emulating a computation as if performed by a trusted third party. Earlier MPC typically require all parties to participate in intensive computations. Recent schemes (Kamara et al., 2011; Scholl et al., 2017; Mohassel and Zhang, 2017) allow these computations to be offloaded to multiple powerful servers without sacrificing original security guarantees, known as the server-aided MPC. In this approach, each party, say $P_{i}$, secretly shares their inputs, $x_{i}$ to servers. These servers then jointly evaluate an MPC that functionally replicates the logic of first reconstructing the secrets $x_{1},x_{2},\ldots$ and then evaluating $f(x_{1},x_{2},\ldots)$. Depending on designs, the result may be directly returned to the parties or re-shared among the servers for later retrieval. By default, SPECIAL considers server-aided MPC with two non-colluding servers.
Secret sharing and secure array. SPECIAL uses the $2$-out-of-$2$ boolean secret share (Araki et al., 2016) over ring $\mathbb{Z}_{2^{32}}$ for securely outsourcing owners’ data and storing query execution results. Specifically, each data, $x$, is divided into two shares: $x1,x2$ that are uniformly distributed over the ring $\mathbb{Z}_{2^{32}}$ such that $x=x1\oplus x2$. Each server $S_{i}$ receives one secret shares, $s_{i}$, where $i\in\{0,1\}$. By retrieving shares from any two servers, an authorized party can successfully reconstruct the value of $x$. However, a single server alone learns nothing about $x$. For clarity and to abstract out the lower-level details, we leverage a logically unified data structure, namely the secure array (Wang et al., 2022; Bater et al., 2018), denoted as $\langle\mathbf{x}\rangle=\left(\langle x_{1}\rangle,\langle x_{2}\rangle,...\right)$, which is a collection of secret-shared relational tuples.
Oblivious (relational) operators. Oblivious operators are specific MPC protocols, applied over secure arrays, that implement database operator functionalities such as filter, join, and aggregation. Specifically, an oblivious operator $\langle\mathbf{o}\rangle\leftarrow\mathsf{\textbf{op}}(\langle\mathbf{a}\rangle,\langle\mathbf{b}\rangle,...)$ accepts one or more secure arrays as inputs, performs database operations over the inputs, and then writes the result to an output secure array. Moreover, the oblivious property stipulates that absolutely no memory access pattern information is disclosed (Chang et al., 2022; Ramachandran and Shi, 2021; Liagouris et al., 2023) (or in other words, the memory access is fully independent from input data) during the protocol’s execution. Since revealing output size can also enable numerous reconstruction attacks, recent designs of oblivious operators also consider hiding output volumes, i.e., by padding the outputs to their maximum sizes. Thus, unless otherwise stated, we assume that all the subsequent oblivious operators mentioned have considered volume-hiding.
Differential privacy (Dwork et al., 2014). DP ensures that modifying a single input data tuple to a mechanism produces only a negligible change in its output. To elaborate, consider $D$ and $D^{\prime}$ as two databases differing by just one tuple—termed as neighboring databases. Then:
Definition 2.1 ($(\epsilon,\delta)$-differential privacy).
Given $\epsilon>0$, and $\delta\in(0,1)$. A randomized mechanism $\mathcal{M}$ is said to be $(\epsilon,\delta)$-DP if for all pairs of neighboring databases $D,D^{\prime}$, and any possible output $o\subset Range(\mathcal{M})$, the following holds:
$\textup{Pr}\left[\mathcal{M}(D)\in o\right]\leq e^{\epsilon}\textup{Pr}\left[\mathcal{M}(D^{\prime})\in o\right]+\delta$
3. Privacy Model
In this section, we formally define the privacy definitions of SPECIAL and explain the corresponding semantics. Similar to previous work on DPSCA (Wang et al., 2022; Zhang et al., 2023), we follow computational DP (SIM-CDP) (Mironov et al., 2009) framework for privacy model formulation.
Definition 3.1 (Secure protocol with DP leakage).
Given a party (owner) $P$ with private data $D$ and a randomized mechanism $\mathsf{Lkg}(D)$, referred to as the leakage profile of $P$’s data. Consider $\Pi$ a secure query processing protocol applied to $D$. The protocol $\Pi$ is said to be secure with DP leakage if, for any honest party $P$ with data $D$ and any probabilistic polynomial time (p.p.t.) adversary $\mathcal{A}$:
  • $\mathsf{Lkg}(D)$ satisfies $(\epsilon,\delta)$-DP (definition 2.1).
  • There exists a p.p.t. simulator $\mathcal{S}$ with only access to public parameters $\mathsf{pp}$ and the output of $\mathsf{Lkg}(D)$ that satisfies:
    (1) $\begin{split}&\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\Pi}(\mathcal{D},\mathsf{pp})=1\right)\right]\\ &\leq\textup{Pr}\left[\mathcal{A}\left(\mathsf{VIEW}^{\mathcal{S}}(F(\mathcal{D}),\mathsf{pp})\right)=1\right]+\mathsf{negl}(\kappa)\end{split}$
where $\mathsf{VIEW}^{\Pi}$, and $\mathsf{VIEW}^{\mathcal{S}}$ denotes the adversary’s view against the protocol execution and the simulator’s outputs, respectively. And $\mathsf{negl}(\kappa)$ is a negligible function related to a security parameter $\kappa$.
Simply put, Definition 3.1 specifies that as long as there is at least one honest owner, the knowledge any p.p.t. adversary can obtain about this owner’s data from observing $\Pi$ execution transcripts is bounded to what can be inferred from the outputs of an $(\epsilon,\delta)$-DP mechanism $\mathsf{Lkg}$. By default, we define $\mathsf{Lkg}$ within $(\epsilon,\delta)$-tuple level DP (Dwork et al., 2010), that’s said primary privacy objective is to protect individual tuples. However, due to group-privacy properties (Kifer and Machanavajjhala, 2011; Xiao and Xiong, 2015) of DP, privacy can extend across multiple tuples and to user-level, as long as a user owns a limited number of tuples. Throughout this paper, our focus remains on developing algorithms that guarantee tuple-level privacy, while upholding all stated privacy guarantees, albeit with potentially different privacy parameters.
Theorem 3.2.
The privacy of SPECIAL adheres to Definition 3.1
Proof.
Due to space concerns, we defer the complete privacy proof of SPECIAL to Appendix A. ∎
4. SPECIAL Synopses
SPECIAL maintains independent DP synopses for all outsourced data. Each synopsis holds noisy statistics that reflect the corresponding data’s distributional traits. As per our core design idea, these statistics will be used to dynamically optimize query processing—for example, by estimating tight compaction bounds or enabling fast access to eliminate unnecessary tuples, which significantly compacts intermediate query sizes, all without any extra privacy cost.
Challenges. A major design challenge is selecting proper synopses to achieve our design goals. This involves decisions such as attribute selection, combinations thereof, and the types of statistics used (e.g., histograms, min/max/avg, frequency counts, etc.). Using too many attributes or leveraging complex (high-dimensional) combinations can introduce significant noise (Zhang et al., 2021) and increase storage costs. Another challenge involves privately releasing synopses to aid in lossless processing, such as result compaction and private indexing. Ensuring lossless compaction requires consistently overestimating the output sizes of subplans. Supporting lossless indexes is even more complex: for a given indexed range $D[\mathsf{lo},\mathsf{hi}]$, we need to overestimate the upper index $\mathsf{hi}$ and underestimate the lower index $\mathsf{lo}$. Unfortunately, traditional DP methods, which typically use symmetric additive noises, struggle to ensure these conditions.
Key ideas. We observe that joins are more resource-intensive than any other operations. For instance a $k$-way oblivious join can incur a $O(n^{k})$ complexity if without optimisations (Bater et al., 2017, 2018; Wang et al., 2022; Zheng et al., 2017; Eskandarian and Zaharia, 2017). Hence, joins require prioritized acceleration. Moreover, synopses available for frequently queried filter attributes on base relations are vital (Wang et al., 2022) as well. These statistics not only help to compact input data at an early stage before heavy joins, but also enable fast indexing (§ 5.2) for quick data access. Hence, to address the first challenge, our key idea is to focus on commonly queried join and base relation filter attributes, as well as their low-dimensional combinations.
To tackle the second challenge, our key idea is twofold. First, to ensure lossless result compaction of filters and indexing, we employ one-sided DP noise to generate private histograms that either overestimate or underestimate attribute distributions consistently. Second, to achieve lossless compaction of join outputs, we integrate noisy max frequency moments (MF) into the synopses. This integration enables us to extend upon advanced join upper bound techniques (Hertzschuch et al., 2021) to privately estimate join compaction sizes without losing data (§ 5.2).
4.1. Synopses generation
Steered by our key ideas, we now elaborate on our synopsis generation details. The general process is outlined as follows: Initially, the SPECIAL servers selects appropriate attributes to generate private synopses, which are then distributed to owners. The owners create the synopses using a DP mechanism and upload them to SPECIAL servers. Finally, the server applies post-processing to these synopses to aid in query processing.
Attributes selection (servers). The first step is to identify a smaller set of attributes for deriving synopses. In general, we consider the existence of a representative workload, $Q_{\mathsf{R}}$ (Kotsogiannis et al., 2019), which can be sourced from a warm-up run of SPECIAL or annotated by the administrator. Note that this process do not incur any private data thus is leakage-free. The servers first identify representative attribute pairs, $\mathsf{pair}=\{\mathsf{pair}_{k}\}_{k\geq 1}$, for each private relation $D\in\mathcal{D}$ via $Q_{\mathsf{R}}$. The designated pairs include: (i) 2-way attribute pairs, which correspond to frequently queried filter-join key combinations; (ii) frequently queried individual attributes not covered by these pairs. By default, each $\mathsf{pair}_{k}=(A_{\mathsf{ft}},A_{\mathsf{j}})$ contains two valid attributes (case i), but either $A_{\mathsf{ft}}$ or $A_{\mathsf{j}}$ may be empty (case ii).
Synopses release (owners). Next, servers pushes the identified pairs to each owner, and subsequently, the owner independently dispatches a private synopsis and returns it to servers. In what follows, we focus on the DP synopses generation mechanism evaluated by each owner, Algorithm 1 illustrates the overall flow.
Algorithm 1 DP synopsis generation $\mathcal{M}_{\mathtt{synop}}$ (in the view of $P$)
1:Input: $\mathsf{pair}=\{\mathsf{pair}_{k}\}_{k\geq 1}$ from servers; private data $D$.
2:$P$ self-determines privacy parameters $\epsilon,\delta$, and init $\mathtt{synop}\leftarrow\emptyset$
3:for each $\mathsf{pair}_{k}$ do
4:     $\mathbf{h}(\mathsf{pair}_{k},D)\leftarrow\mathsf{HistGen}(\mathsf{pair}_{k},D)$
5:DP histograms:
6:     $\mathbf{h}^{+}(\mathsf{pair}_{k},D)\leftarrow\mathbf{h}+\mathsf{Lap}^{+}(\epsilon,\delta,\mathbf{h}.\mathsf{shape})$
7:     $\mathbf{h}^{-}(\mathsf{pair}_{k},D)\leftarrow\mathbf{h}+\mathsf{Lap}^{-}(\epsilon,\delta,\mathbf{h}.\mathsf{shape})$
8$\triangleright$ adding independently sampled noise to every bin of $\mathbf{h}^{+}$, $\mathbf{h}^{-}$
9:DP max frequencies:
10:     if $A_{\mathsf{j}}\in\mathsf{pair}_{k}=\emptyset$ or $A_{\mathsf{j}}$ is unique valued then   $\mathtt{MF}_{k}=\emptyset$
11:     else if $A_{\mathsf{ft}}\in\mathsf{pair}_{k}\neq\emptyset$ then
12$\triangleright$ assuming $\mathbf{h}$ partitions $dom(A_{\mathsf{ft}})$ into $\{B_{1},...,B_{m}\}$
13:         $D^{\ell}\leftarrow\sigma_{A_{\mathsf{ft}}\in B_{\ell}}(D)$ for $\ell=1,2,...,m$
14:         compute noisy MF table, $\mathtt{MF}_{k}=\{\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})\}_{1\leq\ell\leq m}$
15:     else $\mathtt{MF}_{k}=\widehat{\mathtt{mf}}(A_{\mathsf{j}},D)$
16:     $\mathtt{synop}\leftarrow\mathtt{synop}\cup(\mathsf{pair}_{k},\{{\bf h}^{+},{\bf h}^{-}\},\mathtt{MF}_{k})$
17:release $\mathtt{synop}$, $\epsilon$, $\delta$ to servers
Generally, we expect owners to set a desired privacy budget for their data (using parameters $\epsilon$ and $\delta$) before initiating the synopsis generation. For each $\mathsf{pair}_{k}$, the owner first creates a histogram $\mathbf{h}(\mathsf{pair}_{k},D)$. By default, we assume there exist global parameters, i.e. bin sizes of each attribute, which guide all owners to partition attributes consistently. Next, the owner derives two noisy histograms, $\mathbf{h}^{+}(\mathsf{pair}_{k},D)$, and $\mathbf{h}^{-}(\mathsf{pair}_{k},D)$ by adding independently sampled one-sided Laplace noises to every bin of $\mathbf{h}(\mathsf{pair}_{k},D)$.
Definition 4.1 (One-sided Laplace variable).
$\mathsf{Lap}^{+}(\epsilon,\delta)=\max(0,z)$ (resp. $\mathsf{Lap}^{-}(\epsilon,\delta)=\min(0,z)$) is a one-sided Laplace random variable in the range of $[0,\infty)$ (resp. $(-\infty,0]$) if $z$ is drawn from a distribution with the following density function
(2) $\begin{split}\textup{Pr}\left[z=x\right]=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon|x-\mu|}\end{split}$
where $\mu=1-\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))$ (resp. $\mu=\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))-1$).
The one-sided Laplace noise can be strictly positive ($\mathsf{Lap}^{+}$) or negative ($\mathsf{Lap}^{-}$). This ensures that $\mathbf{h}^{+}$ always overestimates the actual histogram, while $\mathbf{h}^{-}$ consistently underestimates it. For the ease of description, we refer $\mathbf{H}(\mathsf{pair}_{k},D)=\{\mathbf{h}^{+},\mathbf{h}^{-}\}$ as the bounding histogram of attribute $\mathsf{pair}_{k}$, with $\mathbf{h}^{+}$, and $\mathbf{h}^{-}$ to be the upper and lower histograms, respectively.
We continue with the details of the (noisy) MFs generation. Assume both $A_{\mathsf{ft}},A_{\mathsf{j}}\neq\emptyset$, and the $\mathbf{h}(\mathsf{pair}_{k},D)$ partitions $dom(A_{\mathsf{ft}})$ into $\{B_{1},...,B_{m}\}$. The mechanism $\mathcal{M}_{\mathtt{synop}}$ generates a table of noisy MFs, $\{\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})\}_{\ell\geq 1}$, where each entry $\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})$ is an independent MF statistic for $A_{\mathsf{j}}$ over specific filtered data on $A_{\mathsf{ft}}$, such that
(3) $\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell})\leftarrow\widehat{\max}_{\epsilon}\left(\mathcal{G}_{\mathbf{count}(A_{\mathsf{j}})}\left(\sigma_{A_{\mathsf{ft}}\in B_{\ell}}(D)\right)\right)$
here $\mathcal{G}_{\mathbf{count}(A_{\mathsf{j}})}$ is a group-by-count operation over $A_{\mathsf{j}}$, and $\widehat{\max}_{\epsilon}$ is a report noisy max mechanism (Dwork et al., 2014). It first adds i.i.d. noise from the exponential distribution $\mathsf{Exp}(\frac{2}{\epsilon})$ to each grouped count, then outputs the largest noisy count. We stress that $\mathcal{M}_{\mathtt{synop}}$ will not generate noisy MFs for non-join key attributes, and when $A_{\mathsf{ft}}$ is empty, a global MF will be generated instead of MF tables (Alg 1:10). Moreover, since SPECIAL enables owners to label attributes as unique-valued, if $A_{\mathsf{j}}$ is known to be unique-valued, then $\smash{\widehat{\mathtt{mf}}(A_{\mathsf{j}},\cdot)}$ is always 1. Nevertheless, as exponential noises are non-negative, thus $\widehat{\mathtt{mf}}\geq\mathtt{mf}$ holds for all cases. Given the discussion so far, we summarize SPECIAL synopses as:
Definition 4.2 (SPECIAL synopses).
Given representative workload $Q_{\mathsf{R}}$, we consider for each relation $D$, its corresponding synopsis $\mathtt{synop}$ to be the collection of $\{(\mathsf{pair}_{k},\mathbf{H}(\mathsf{pair}_{k},D),\mathtt{MF}_{k})\}_{k\geq 1}$, such that
  • $\mathsf{pair}_{k}$ is a frequently queried attribute (pair) over $D$ found in $Q_{\mathsf{R}}$.
  • $\mathbf{H}$ is the private bounding histogram for $\mathsf{pair}_{k}$ over $D$.
  • $\mathtt{MF}_{k}$ represents a collection of privately overestimated join key MFs that corresponds to a group of tuples in $\mathbf{h}$, categorized by $\mathsf{pair}_{k}.A_{\mathsf{ft}}$.
Theorem 4.3.
Given $|\mathsf{pair}|=c$, $\epsilon,\delta>0$, the synopsis generation (Algorithm 1) is $(\hat{\epsilon},\hat{\delta})$-DP where $\hat{\epsilon}\leq 6\epsilon\sqrt{c\ln(1/\delta)}$, and $\hat{\delta}=(c+1)\delta$
For space concern, we move the complete proof to Appendix A. In a sketch, adding $\mathsf{Lap}^{+}$ (or $\mathsf{Lap}^{-}$) to a single bin is $(\epsilon,\delta)$-DP. By parallel and sequential composition, generating $\mathbf{{H}}$ is $(2\epsilon,2\delta)$-DP. Moreover, each report noisy max is $(\epsilon,0)$-DP, and by parallel composition, the generation of the entire MF table is also $(\epsilon,0)$-DP. In this way, we know that the generation of each $(\mathsf{pair}_{k},\mathbf{H},\mathtt{MF}_{k})$ is at most $(3\epsilon,\delta)$-DP. Given there are in total $c$ such pairs, and thus the total privacy loss is subject to $c$-fold advanced composition (Dwork et al., 2014).
Synopsis transformations. We say that one can perform transformations on released synopses without incurring extra privacy loss, per the post-processing theorem of DP (Dwork et al., 2014). Now, we briefly outline key synopsis transformations relevant to the SPECIAL design. First, given any (2d) bounding histogram $\mathbf{H}(\mathsf{pair},D)$ with both $A_{\mathsf{ft}},A_{\mathsf{j}}\in\mathsf{pair}$ are non-empty, one can derive the (1d) bounding histograms, i.e. $\mathbf{H}(A_{\mathsf{j}},D)$ and $\mathbf{H}(A_{\mathsf{ft}},D)$, for any single attribute $A_{\mathsf{ft}}$ or $A_{\mathsf{j}}$ by marginal sums $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbf{H}(\mathsf{pair},D)$ over $A_{\mathsf{j}}$ or $A_{\mathsf{ft}}$, respectively. This enables the creation of statistics on individual attributes, even when $A_{\mathsf{ft}}$ and $A_{\mathsf{j}}$ are not included as a standalone synopsis attribute. Moreover, it’s possible to derive relevant join key statistics even following a selection on the base relation. For example, given $A_{\mathsf{ft}},A_{\mathsf{j}}\in\mathsf{pair}\neq\emptyset$, and let $D^{\prime}\leftarrow\sigma_{A_{\mathsf{ft}}\in vals}(D)$, one can obtain the (1d) bounding histogram $\mathbf{H}(A_{\mathsf{j}},D^{\prime})$ by conducting a selective marginal sum of $\mathbf{h}^{+},\mathbf{h}^{-}\in\mathbf{H}(\mathsf{pair},D)$ over bins of $A_{\mathsf{ft}}$ that intersect with $vals$. Beyond bounding histograms, join key MFs over pre-filtered data can also be computed by
(4) ${}\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\prime})=\min\left(\textstyle\sum_{B_{\ell}\cap vals\neq\emptyset}\widehat{\mathtt{mf}}(A_{\mathsf{j}},D^{\ell}),\widehat{\mathtt{mf}}(A_{\mathsf{j}},D)\right)$
Note that $\widehat{\mathtt{mf}}(A_{\mathsf{j}},D)$ exists if $A_{\mathsf{j}}$ is also included as a standalone synopsis attribute; otherwise, Eq 4 yields only the first term.
5. SPECIAL Primitives
We now outline key secure database operations offered by SPECIAL. A major challenge in designing synopsis-assisted secure processing is the evident knowledge gap on how private synopses could potentially accelerate oblivious query processing. To bridge this gap, we explore various uses of synopses, such as creating private indexes (SPEidx 5.2) and designing compacted oblivious operations (SPEop 5.3). Moreover, as discussed earlier, joins are the most intensive operations. Hence, we extensively optimize join algorithms, synergizing both private indexing and compaction techniques to develop a novel, parallel-friendly oblivious join (§ 5.3). Another challenge in primitive design arises from ensuring lossless guarantees. To address this, we integrate mechanisms into our designs that pessimistically estimate selection cardinalities, indexing ranges, and complex join sizes, using a combination of available synopses and advanced upper bound techniques (Hertzschuch et al., 2021). For clarity, we’ll assume, WLOG., that all input relations below are of size $n$ and all 1d histograms contain $m$ bins.
5.1. Basic Operations
SPECIAL supports conventional fully-oblivious operators (Bater et al., 2017). In general, these operators function logically the same as non-private ones, but their execution transcripts are completely data independent, with results consistently padded to the worst-case maximum.
Default data access SeqACC. By default, all query executions proceed with a standard data access, loading all data into a secure array using a sequential scan [^2] In the context of a secret-shared array $\langle D\rangle$, this step can be each server sequentially load the respective secret shares of $D$ into memory, readying them for MPC.. Additionally, each loaded tuple gets a secret bit (initially ‘0’), $\mathsf{ret}$, used to mark tuple validity.
SELECT. is the secure version of filter $\sigma_{p}(R)$. It conducts a linear scan over secure array $\langle R\rangle$ for relation $R$ and updates the $\mathsf{ret}$ bit: ‘1’ for tuples satisfying predicate $p$ and ‘0’ for others. Every tuple’s $\mathsf{ret}$ bit is updated, regardless of whether it matches the predicate.
PROJECT. discards irrelevant attributes for relation stored in $\langle R\rangle$. Note that the secret bit $\mathsf{ret}$ cannot be discarded by a PROJECT.
JOIN. implements the secure version of $\theta$-join, $R_{0}\Join_{\theta}R_{1}$. It processes two secure arrays $\langle R_{0}\rangle$ and $\langle R_{1}\rangle$, computing their cartesian product $\langle R_{0}\times R_{1}\rangle$ and then using SELECT to mark the joined tuples. It’s important to note that JOIN expands output, always padding it to the worst-case max (product of input sizes).
COUNT, SUM, MIN/MAX are the aggregation primitives. The operators linearly scan a secure array and continuously update a secret-shared aggregation value after each tuple access.
ORDER-BY, DISTINCT, GROUP-BY(AGG). are built on the oblivious sort primitive (Batcher, 1968). Specifically, ORDER-BY obliviously sorts a secure array according to a given attribute. DISTINCT sorts a secure array first, then performs a linear scan to identify unique tuples. For instance, among consecutively sorted identical tuples, it sets the $\mathsf{ret}$ bit of the last tuple to ‘1’ and the rest to ‘0’. GROUP-BY(AGG) starts by using DISTINCT to identify unique tuples. For each distinct tuple ($\mathsf{ret}$ set to ‘1’), it appends an aggregation value derived from the tuple and a dummy attribute (e.g., ‘-1’) for non-distinct tuples.
5.2. SPECIAL Index SPEidx
In conventional databases, indexes are powerful data structures that map attribute values to positions within a sorted array, enabling fast data access. However, generalizing such feature to SCA remains largely underexplored with existing solutions facing significant limitations such as high data loss (Roy Chowdhury et al., 2020), dependence on complex data structures filled with dummy tuples (Zhang et al., 2023; Bogatov et al., 2021), and limited query capabilities (i.e., linear queries across fixed attributes). Moreover, these methods only index base relations. SPEidx innovatively overcomes these issues by enabling the creation of lossless indexes directly on outsourced data or query intermediate results, without the need for additional structures or dummy injections.
In general, SPEidx builds upon the typical indexing model that utilizes cumulative frequencies (CF) (Kraska et al., 2018). Specifically, given $D$ sorted by $A\in attr(D)$, all records $t\in D$ where $t.A=x$ can be indexed by the interval $[g(x-1),g(x)]$, where $g(x)=|\{t\mid t.A\leq x\}|$ is the CF function. For better illustration, we show an example index lookup in Figure 2: to get all records with an attribute value of 24, one may compute $[g(23),g(24)]=[217,248]$ and access the relevant data from the subset $D[217:248]$. To make this indexing method private and lossless, the key idea of SPEidx is to derive two noisy CF curves from synopses (e.g. bounding histograms). One curve, $g^{+}(x)$, consistently overestimates $g(x)$, while the other, $g^{-}(x)$, consistently underestimates it. Then for any attribute value $x$, we can now derive a private interval $[g^{-}(x-1),g^{+}(x)]$ that losslessly indexes all desired records. For instance, as illustrated in Figure 2, the SPECIAL index might estimate the index range for attribute value 24 as $[g^{-}(23),g^{+}(24)]=[198,267]$.
Figure 2. True index (left) vs. SPECIAL index (right) for $x=24$
In what follows, we provide the formal explanations on how SPEidx derives indexes from DP synopses. Specifically, SPEidx first determines the bounding histograms $\mathbf{H}(A,D)$, which may be either transformed from an available 2D histogram $\mathbf{H}(\mathsf{pair},D)$ with $A\in\mathsf{pair}$, or sourced directly if $\mathbf{H}(A,D)$ is already included in the synopses. It then constructs the noisy mapping as follows:
Definition 5.1 (SPECIAL index).
Given $D$ sorted by $A$, the bounding histogram $\mathbf{H}(A,D)=\{\mathbf{h}^{+},\mathbf{h}^{-}\}$, and assume $\mathbf{h}^{+}=(c^{+}_{1},...,c^{+}_{m})$, $\mathbf{h}^{-}=(c^{-}_{1},...,c^{-}_{m})$ partitions $dom(A)$ into $\{B_{1},...,B_{m}\}$. We say $\textsc{SPEidx}(A,D)=\{\mathsf{idx}_{i}=[\mathsf{lo}_{i},\mathsf{hi}_{i}]\}_{1\leq i\leq m}$ is the SPECIAL index of $D$ over $A$ with:
  • $\forall~{}i\geq 1$, $\mathsf{hi}_{i}=\min(|D|,\sum^{i}_{k=1}c^{+}_{k})$.
  • $\mathsf{lo}_{1}=0$, and $\forall~{}i\geq 2$, $\mathsf{lo}_{i}=\sum^{i-1}_{k=1}\max(0,c^{-}_{k})$.
By this construction, all tuples $t\in D$ such that $t.A\in B_{i}$ will be organized into the subset $D[\mathsf{idx}_{i}]\subseteq D$. This subset can be quickly accessed if $D$ is already sorted, without the need for special data structures or inclusion of dummy tuples. Depending on how bounding histograms are constructed, $\textsc{SPEidx}(A,D)$ can support indexing lookups with varying granularity. This can range from indexing individual attribute values (where each $B_{i}$ corresponds to a single domain value) to indexing a range of of values. The bounding histogram’s pessimistic estimation ensures that all tuples where $t.A\in B_{i}$ are accurately contained within $D[\mathsf{idx}_{i}]$, thereby achieving lossless indexing. In contrast to existing methods that are limited to indexing base relations (Roy Chowdhury et al., 2020; Bogatov et al., 2021; Zhang et al., 2023), SPEidx extends its capabilities to create private indexes on query intermediate results. For instance, consider $D^{\prime}\leftarrow\sigma_{A^{*}\in vals}(D)$ where the attribute pair $(A^{*},A)$ is included in $\mathtt{synop}$. Here, SPEidx can derive $\mathbf{H}(A,D^{\prime})$ from $\mathbf{H}(A^{*},D)$ and subsequently build indexes on $D^{\prime}$. Importantly, since index creation is a post-processing activity using available DP synopses, it incurs no additional privacy loss.
Indexed store and fast data access IdxAcc. SPEidx enables the exploration of a new storage layout for organizing outsourced data, namely indexed datastore, that facilities private data access. Specifically, by analyzing a representative workload $Q_{\mathsf{R}}$, one may identify the “hottest” attribute per base relation, sort them according to the “hottest” attribute, and then build indexes over the sorted data. This storage layout enables fast indexed access (IdxAcc) to retrieve a compact subset of data from the outsourced relations, thereby eliminating the need for a full table sequential scan (SeqAcc) potentially bypassing selection operations (§ 5.3). We emphasize that the SPECIAL design does not require replicating the outsourced datastore to accommodate multiple query types (Bogatov et al., 2021; Zhang et al., 2023). However, creating compact replicas (e.g., column replicas (Huang et al., 2020) over frequently queried attributes) can be optionally employed to enhance query processing speed. Moreover, the generation of all aforementioned objects (indexed store and column replicas) requires only three primitives: projection, oblivious sorting, and SPEidx, all achieved without extra privacy loss. In other words, this implies that one can selectively adjust these objects to align with dynamic query workloads, without incurring any privacy costs.
5.3. SPECIAL Operators SPEop
We introduce SPEop, a set of novel synopsis-assisted operators that maintain full obliviousness (e.g., operators’ execution cause no privacy loss) while enabling lossless compaction. To our knowledge, SPEop is the first primitive of its kind in any SCA system.
Oblivious compaction: OPAC. We first introduce the fundamental operation of oblivious compaction, a key component for many SPEop primitives. Given input $\langle R\rangle$, OPAC sorts it based on the secret bit $\mathsf{ret}$, moving tuples with $\mathsf{ret}=$ ‘1’ to the front. Then, OPAC retains only the first $k$ tuples from the sorted array. The compaction is lossless if $k$ is greater than or equal to the number of tuples with $\mathsf{ret}=$ ‘1’; otherwise, it is lossy.
SPECIAL selections: (OP)SELECT, (SP)SELECT, (DC)SELECT. Let $R$ to be a relation and $A\in attr(R)$, we now introduce three advanced selections that implements $\sigma_{A\in vals}(R)$.
Algorithm 2 $\mathtt{CardEst}(\sigma_{A\in vals}(R),\mathtt{synop})$
1:$\mathbf{h}=\emptyset$, $c=0$
2:if $\mathbf{H}(A,R)\in\mathtt{synop}$ then $\mathbf{h}\leftarrow\mathbf{h}^{+}\in\mathbf{H}(A,R)$
3:else if $\exists~{}\mathsf{pair}\in\mathtt{synop}$, s.t. $A\in\mathsf{pair}$ then
4:     $\mathbf{h}\leftarrow$ marginal sum $\mathbf{h}^{+}\in\mathbf{H}(A,R)$ over $\left(\mathsf{pair}\setminus A\right)$.
5:else return $cs=|R|$
6:return $cs=\min(|R|,\sum_{i=1}^{m}c_{i}\in\mathbf{h}:\left(B_{m}\cap vals\neq\emptyset\right))$
(OP)SELECT. is mainly implemented based on the oblivious compaction (OPAC) operation. Specifically, the operation first conducts a standard SELECT on the input secure array $\langle R\rangle$ to label selected tuples, followed by an OPAC to to eliminate a large portion of non-matching tuples. To determine the compaction size $cs$, (OP)SELECT examines the synopsis of $R$ and pessimistically estimates the cardinality of $\sigma_{A\in vals}(R)$ as shown in Algorithm 2. Since $cs$ never underestimates the actual cardinality, and thus, the compaction is lossless with no missing tuples. Moreover, as OPAC is fully oblivious and $cs$ is determined completely from post-processing over DP synopsis, thus, (OP)SELECT causes no privacy loss.
(SP)SELECT. The running complexity of (OP)SELECT depends on OPAC, which is typically linearithmic (see § 6.1 or (Sasy et al., 2022)). However, when $\mathtt{CardEst}(\sigma_{A\in vals}(R),\mathtt{synop})$ is relatively small, oblivious selection can be achieved without necessarily incurring linearithmic cost. Specifically, we consider (SP)SELECT, which first creates an empty output array $\langle R_{o}\rangle$ with size equals to $cs$ before any computations. Next, it evaluates two linear scans over $\langle R\rangle$, where the first scan obliviously marks all selected tuples, and in the second scan, it privately writes all marked tuples into $\langle R_{o}\rangle$. Specifically, in the second scan, (SP)SELECT internally maintains the last actual write position $\mathsf{idx}$ in $\langle R_{o}\rangle$. Then for every newly accessed tuple $\langle t\rangle$ in $\langle D\rangle$, a write action occurs on all tuples in $\langle R_{o}\rangle$. If $\langle t\rangle$ is selected, then an actual write is made that writes $\langle t\rangle$ to $\langle R_{o}[\mathsf{idx}+1]\rangle$ and a dummy write [^3] In the context of the secret-shared secure array $\langle\mathbf{a}\rangle$, a dummy write to $\langle\mathbf{a}[i]\rangle$ is simply a re-sharing of $\mathbf{a}[i]$ through secure protocols without changing its value. is made to elsewhere. If not, dummy writes are made throughout $\langle R_{o}\rangle$. Similarly, (SP)SELECT does not yield additional privacy loss, and the complexity is linear in its input.
(DC)SELECT. Finally, if the underlying data is already indexable on $A$, a direct selection over the data can be applied that bypasses secure computations. The operator simply looks up $\textsc{SPEidx}(A,R)$, and accesses $R[a,b]$, where $a=\min_{i}(\mathsf{idx}_{i}.\mathsf{lo}),b=\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})$, where $\mathsf{idx}_{i}$ dentoes the index in $\textsc{SPEidx}(A,R)$ where $B_{i}\cap vals\neq\emptyset$.
SPECIAL join: (MX)JOIN. We now introduce a novel MF-Index based oblivious join operation. The advancements of (MX)JOIN stand out in two key aspects compared to existing oblivious joins. First, compared to the standard JOIN, (MX)JOIN stands out for its ability to significantly compact the output size, coupled with a highly parallelizable fast processing mode. Second, existing oblivious joins with DP leakages typically require additional privacy budget to learn join sensitivity (Dong et al., 2022) or necessitate truncations on joined tuples (Wang et al., 2022; Bater et al., 2018). Moreover, (MX)JOIN is unique as the first oblivious join that enables lossless output compaction without extra privacy loss. We illustrate the construction details in Algorithm 3.
Algorithm 3 (MX)JOIN (base and pre-filtered relations)
1:Input: relations $R_{0}$, $R_{1}$; join attribute $A_{\mathsf{j}}$.
2:if $\mathsf{MXReady}(R_{0},R_{1})==\mathsf{True}$  then  $\mathsf{BucketJoin}(R_{0},R_{1},A_{\mathsf{j}})$
3:else if $R_{0}$, $R_{1}$ are either base or pre-filtered relation then
4:     for $b\in\{0,1\}$ do
5:         derive $\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{b})$ from $\mathtt{synop}_{b}$ (§ 4)
6:         build index $\textsc{SPEidx}(A_{\mathsf{j}},R_{b})=\{\mathsf{idx}_{i}\}_{i=1,..,m}$ (§ 5.2)
7:     if $\forall b~{}$, $\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{b})$, and $\textsc{SPEidx}(A_{\mathsf{j}},R_{b})\neq\mathsf{null}$  then
8:         oblivious sort $R_{0}$, $R_{1}$ on $A_{\mathsf{j}}$, $\mathsf{BucketJoin}(R_{0},R_{1},A_{\mathsf{j}})$
9:     else assert “not applicable for (MX)JOIN”
10:$\mathsf{BucketJoin}(R_{0},R_{1},A_{\mathsf{j}})$:
11:for $i=1,2,...,m$ do
12:     compute $O_{i}\leftarrow(R_{0}[\mathsf{idx}_{i}]\Join_{A_{\mathsf{j}}}R_{1}[\mathsf{idx}_{i}])$ via basic JOIN
13:     $cs_{i}\leftarrow\min\left(\frac{|R_{0}[\mathsf{idx}_{i}]|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})},\frac{|R_{1}[\mathsf{idx}_{i}]|}{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}\right)\times{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})}\cdot{\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})}$
14:     $R_{\mathsf{out}}\leftarrow R_{\mathsf{out}}\cup{\texttt{OPAC}}(O_{i},cs_{i})$
15:return $R_{\mathsf{out}}$
In general, (MX)JOIN can be applied to two types of data: the base and pre-filtered relations where the join key attribute is contained by the synopses. Specifically, (MX)JOIN starts with computing the join key MFs (Alg 3:4) and constructing private indexes (Alg 3:5) for both inputs. All these operations are conducted through “privacy cost-free” transformations using available DP synopses. Once these objects are obtained, the algorithm employs oblivious sort to rearrange both inputs (Alg 3:6,7), rendering them indexable with tuples logically distributed into independent buckets by join key values. Next, (MX)JOIN simply adopts standard JOIN to join tuples exclusively within the same buckets (Alg 3:10). Finally, (MX)JOIN performs per-bucket output compaction, where it first determines the MF join bound (Hertzschuch et al., 2021) for each bucket join and invokes OPAC to compact the output according to the learned size (Alg 3:11,12). As bucket-wise operations are independent, the aforementioned steps lend themselves well to parallelized processing. For better illustration, we visualize (MX)JOIN’s processing flow in Figure 3.
Figure 3. Example processing flow of (MX)JOIN.
Since (MX)JOIN derives join compaction sizes completely from post-processing of DP synopses, it thus incurs no extra privacy loss. Moreover this design choice also eliminates the need for computing join sensitivities — a major hurdle in existing — as accurate join sensitivity estimation is hard and typically requires a significant privacy budget (Dong et al., 2022). Additionally, the noisy MF bounds guarantee that compaction sizes are consistently overestimated, ensuring lossless compaction of join results.
6. SPECIAL Planner
A major limitation of existing DPSCA designs is sub-optimal execution planning. Without the ability to pre-estimate intermediate result sizes, traditional methods cannot identify the most efficient execution strategies. To address this, SPECIAL employs an innovative query planner that leverages available synopses for plan size estimation, enabling optimal execution planning.
At a high level, our planner is modeled on the Selinger-style (cost-based) optimizer (Blasgen et al., 1981). It uses a bottom-up, dynamic programming approach to enumerate all equivalent secure execution plans for a given query and estimate their costs to select the optimal one. However, challenges in the design space still remain. First, the introduction of SPECIAL primitives can significantly impact the cost modeling for oblivious operations, rendering existing models (Bater et al., 2018; Liagouris et al., 2023) ineffective. Additionally, the equivalent plan search space is extensive, requiring further strategies to simplify the search space. To address these challenges, we first conduct a comprehensive analysis of the complexities of SPECIAL primitives and develop a new cost model (§ 6.1) for SPECIAL planner. Next, we design protocol-specific heuristics (§ 6.2) tailored for the SPECIAL planner to simplify the plan search space. Note that existing features of the Selinger-style optimizer, such as plan enumerations, are not within our contributions. For detailed understanding of these features, we encourage readers to explore the original IBM System R paper (Blasgen et al., 1981).
6.1. Cost Model
Similar to the cost model proposed by Shrinkwrap (Bater et al., 2018), we view the cost of a secure execution plan cost as a combination of each operator’s I/O and secure evaluation costs. In addition, we formulate these costs as functions over their processed data sizes. Specifically, given a plan with $\ell$ operators, $\mathsf{op}_{1},...,\mathsf{op}_{\ell}$, and let $\mathbf{I}=\{{I}_{1},...,{I}_{\ell}\}$, $\mathbf{O}=\{{O}_{1},...,{O}_{\ell}\}$, to be the input and output sizes of each operators. The plan cost is formulated as:
(5) $\mathsf{Cost}=\textstyle\sum_{i=1}^{\ell}C^{\mathsf{op}_{i}}_{\mathsf{in}}(I_{i})+C_{\mathsf{eval}}^{\mathsf{op}_{i}}(I_{i})+C^{\mathsf{op}_{i}}_{\mathsf{out}}(O_{i})$
Here, $C_{\mathsf{in}}$ represents the data access cost (input I/O), primarily capturing the expenses when moving data from persistent storage to an in-memory secure array. If the accessed data has already loaded into memory, $C_{\mathsf{in}}$ becomes negligible. $C_{\mathsf{out}}$ denotes the output I/O cost, modeling the expenses when writing operator results into output arrays. $C_{\mathsf{eval}}$ accounts for the secure computing expenses for evaluating each operators, typically constituting the dominant cost. Note that, in practice, the exact formulas for $C_{\mathsf{in}},C_{\mathsf{out}}$, and $C_{\mathsf{eval}}$ can vary depending on the specific secure protocol employed (garbled circuits (Yao, 1986), secret sharing (Micali et al., 1987), etc.) as well as the particular hardware configurations in use. Nonetheless, the understanding of the asymptotic costs is adequate for comprehending the principles of our query planning and optimization strategies (Liagouris et al., 2023; Bater et al., 2018). In what follows, we provide detailed analysis on the asymptotic costs for each SPECIAL operator. Similarly, we assume that all input data sizes mentioned henceforth in this section are of size $n$, and all 1D histograms have $m$ bins. Table 1 summarizes the operator costs.
Table 1. Asymptotic costs for secure operators
Operator Input I/O ($C_{\mathsf{in}}$) Eval. ($C_{\mathsf{eval}}$) Output I/O ($C_{\mathsf{out}}$)
PROJECT $O(n)$ N/A $O(n)$
Agg. $O(n)$ $O(n)$ $O(1)$
Group & Order $O(n)$ $O(n\log^{2}n)$ $O(n)$
SELECT $O(n)$ $O(n)$ $O(n)$
(OP)SELECT $O(n)$ $O(n\log n)$ $\mathsf{hist\_bound}$
(SP)SELECT $O(n)$ $O(n)$ $O(1)$
(DC)SELECT $\mathsf{idx\_bound}$ N/A N/A
JOIN $O(n)$ $O(n^{2})$ $O(n^{2})$
(MX)JOIN $O(n)$ $O(n^{2})^{*}$ $\mathsf{mf\_bound}$
Assuming the maximum size of the indexed buckets is bounded by $O(\frac{n}{\log n})$.
Oblivious sorting and compaction. Oblivious sorting and compaction (OPAC) are crucial for many SPEop operators, so we will begin our analysis with these two primitives. While oblivious sorting algorithms with optimal $O(n\log n)$ complexity exist, they often necessitate either impractically large constants (Ajtai et al., 1983; Goodrich, 2014) or client-side memory (Asharov et al., 2020), both do not fit with SCA scenario. Consequently, we will consider the well-established bitonic sorting based implementation for oblivious sort, which come with $O(n\log^{2}n)$ complexity. Nonetheless, efficient OPAC implementations with $O(n\log n)$ complexity remain achievable (Sasy et al., 2022)
Projection, grouping and aggregation. The PROJECT accesses private relations and discards unnecessary columns independently on each server, without secure computations. Thus, I/O costs dominate this operation, with both input and output costs bounded by $O(n)$. The costs of ORDER-BY, DISTINCT, and GROUP-BY are primarily due to oblivious sorting, resulting in a complexity of $O(n\log^{2}n)$. Additionally, as these operators do not reduce output sizes, both $C_{\mathsf{in}}$ and $C_{\mathsf{out}}$ are bounded by $O(n)$. Finally, the cost of aggregations, i.e. COUNT, SUM, and MIN/MAX subjects to a oblivious linear scan, typically outputting a single secret-shared value. Hence, its $C_{\mathsf{eval}}$ is bounded by $O(n)$, with $C_{\mathsf{in}}$ at $O(n)$ and $C_{\mathsf{out}}$ at $O(1)$.
Selections. The primary cost of SELECT stems from an oblivious linear scan, making $C_{\mathsf{eval}}$ within $O(n)$. Since SELECT does not shrink the output size, both $C_{\mathsf{in}}$ and $C_{\mathsf{out}}$ are within $O(n)$. (OP)SELECT requires an oblivious compaction (OPAC) before writing outputs, where OPAC usually yields an $O(n\log n)$ complexity (Sasy et al., 2022). Consequently, its $C_{\mathsf{eval}}$ is bounded by $O(n\log n)$ with input I/O cost same as SELECT. However, as (OP)SELECT compacts output size, $C_{\mathsf{out}}$ is reduced to $\mathsf{hist\_bound}=O\left(\sum_{B_{i}\cap vals\neq\emptyset}|B_{i}|\right)$, where $B_{i}\cap vals\neq\emptyset$ are bins in the synopsis histogram intersecting with selection conditions. If $\sum_{B_{i}\cap vals\neq\emptyset}|B_{i}|\sim O(1)$, (SP)SELECT becomes preferable, with its running cost dominated by a two-phase linear scan, and thus $C_{\mathsf{eval}}$ is now $O(n)$, and the output cost is $O(1)$. (DC)SELECT is the most efficient selection, though it requires indexable input data. It simply loads indexed data into secure arrays without secure computations or result write-backs. As such, both $C_{\mathsf{evals}}$ and $C_{\mathsf{out}}$ are negligible, with $C_{\mathsf{in}}$ within $\mathsf{idx\_bound}=O\left(\max_{i}(\mathsf{idx}_{i}.\mathsf{hi})-\min_{i}(\mathsf{idx}_{i}.\mathsf{lo})\right)$. Here, $\mathsf{idx}_{i}$ are indexed regions that intersect with selection conditions.
Joins. Both JOIN and (MX)JOIN have $O(n)$ data access costs, but differ in $C_{\mathsf{eval}}$ and $C_{\mathsf{out}}$. JOIN, using a Cartesian product approach, doesn’t reduce output size, so both $C_{\mathsf{eval}}$ and $C_{\mathsf{out}}$ are within $O(n^{2})$. Compared to JOIN, in the worst-case scenario where the join keys follow a highly biased distribution, i.e. max bucket size reaches $O(n)$, (MX)JOIN’s asymptotic cost is at most $O(n^{2}\log n)$. However, when join keys are distributed more uniformly, the cost can be asymptotically better. For instance, with $m=\log n$ and assuming a max bucket size of $O(\frac{n}{\log n})$, each bucket join costs $O(\frac{n^{2}}{\log n})$, leading to a total cost of $O(n^{2})$, equivalent to JOIN. Recall that bucket joins in (MX)JOIN can be executed concurrently, hence, the processing latency is indeed dominated by the bucket-wise cost, i.e. $O(\frac{n^{2}}{\log n})$. Additionally, the output cost is lowered from $O(n^{2})$ to the sum of per-bucket MF upper bounds (Alg 3:11), which can be substantially less if the join key MFs are relatively low.
6.2. Heuristics
H-1. Filter push down. Filter pushdown is a widely used optimization technique in conventional databases, where it involves moving the selection operation to the earliest stage in the data processing pipeline, such as directly to the base relation. This reduces the amount of data that needs to be loaded, transferred, and processed, by following operators especially complex joins. However, in SCA designs, data obliviousness typically necessitates padding selection sizes to the worst-case scenario, rendering filter pushdown ineffective (Liagouris et al., 2023; Eskandarian and Zaharia, 2017; Zheng et al., 2017; Poddar et al., 2021). With SPEop selections, the selection outputs can be significantly compacted, making filter pushdown effective again. Therefore, we include filter push down as one of the optimization heuristic for SPEplan.
H-2. Predicates fusion. Let $R$ to be any relation, $A_{1},A_{2},...,A_{k}\subseteq attr(D)$, and $\mathbf{v}=\{v_{1},v_{2},...,v_{k}\}$. We say that for multiple selection over $R$ such that $\sigma_{A_{1}\in v_{1}}(...\sigma_{A_{k}\in v_{k}}(R))$, one can always fuse them into one selection $\sigma_{\mathbf{A}\subset\mathbf{v}}(R)$. This can reduce the number of secure computation invocations from $k$ rounds to just one. Additionally, the selection size can be estimated as $\min_{A_{i}}\left(\mathtt{CardEst}(\sigma_{A_{i}\in v_{i}}(R))\right)$.
H-3. Join statistics propagation. A key property of SPECIAL join, (MX)JOIN, is that the output is already indexed and bucketized by the join key. Therefore, for any output $R$ of (MX)JOIN computing $R_{0}\Join_{A_{\mathsf{j}}}R_{1}$, a new index $\textsc{SPEidx}(A_{\mathsf{j}},R)$ across $R$ can be easily derived. Moreover, as per (Hertzschuch et al., 2021), one can also update the MF for $R$ by computing $\widehat{\mathtt{mf}}(A_{\mathsf{j}},R)=\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{0})\times\widehat{\mathtt{mf}}(A_{\mathsf{j}},R_{1})$. As a result, we say that the output of (MX)JOIN as MF-and-index-ready, enabling direct application of another (MX)JOIN on the same join attribute.
7. Evaluation
In this section, we present evaluation results of our proposed framework. Specifically, we address the following questions:
  • Question-1: Does SPECIAL offer efficiency advantages over current SCA approaches? To what extent can SPECIAL enhance efficiency compared to existing state-of-the-arts (SOTAs)?
  • Question-2: For the DP-based SPECIAL design, is there a trade-off between privacy and efficiency? Can we adjust privacy levels to achieve various efficiency objectives?
  • Question-3: Can SPECIAL scale complex analytical (e.g. multi-way join) queries to large-scale (multi-million rows) datasets?
7.1. Experimental setups
We briefly outline our experimental setup, covering the baseline systems, prototype SPECIAL implementation, workloads, datasets, and default configurations.
Baseline systems and SPECIAL prototype. We compare SPECIAL with two baseline systems: Shrinkwrap (Bater et al., 2018), the SOTA DPSCA, and SMCQL (Bater et al., 2017), a typical SCA system using exhaustive padding for leakage hiding, also used as a baseline for Shrinkwrap. For consistency, we consider the same circuit-model implementations for both baseline systems and the SPECIAL prototype. While some works (Poddar et al., 2021; Liagouris et al., 2023) similar to SMCQL use exhaustive padding but improve efficiency through protocol-level optimizations, we exclude them from our benchmarks for fair comparison concerns. However, it’s important to note that these optimizations are orthogonal to both the baseline systems and our designs. We re-implement important query features for the two baseline systems, and built SPECIAL using the same MPC package, specifically the sh2pc-0.2.2 protocol within EMPtoolkit-0.2.5. Since managing DP synopses and generating indexes do not involve secure computations, thus we implement the related features using Python (3.7) scripts. All implementations are open-sourced (Lovingmage, 2024).
Datasets and workloads. Although SMCQL and Shrinkwrap use the HealthLNK dataset, it is not public data and thus not accessible to the broader research community. Hence, we developed a new benchmark based on a publicly available dataset, and mimic Shrinkwrap’s workload to create a set of testing queries. Specifically, we use the Czech Financial Dataset (Rel, [n.d.]), an anonymized Czech bank transaction dataset for loan applications. This dataset comprises 8 relational tables with a total of 55 columns and 1.1 million rows. For scaling experiments, we use the raw data and schema to generate synthetic data with up to 10 million rows. Shrinkwrap’s benchmark evaluates four queries: a range query, two 2 binary joins, and a 3-way join. Mirroring this approach, we designed eight query workloads based on the Financial data. All workloads are selection-projection-join (SPJ) queries followed by statistical aggregations such as count, count distinct, and group by. The workloads range from simple linear queries to complex multi-way joins. Due to space limitations, a brief summary of the workloads is provided in Table 2, with the full specification in Appendix B.
Table 2. Query workloads
Query No. Type Description
Q1 Linear Point query with single predicate.
Q2 Linear Range query with conjunctive predicates.
Q3 Binary Join Non-expanding binary join.
Q4 Binary Join Expanding binary join.
Q5 Multi Join 3-way mixed non-expanding and expanding joins.
Q6 Multi Join 3-way all expanding joins.
Q7 Multi Join 4-way mixed non-expanding and expanding joins.
Q8 Multi Join 5-way mixed non-expanding and expanding joins.
Default configurations. For SPECIAL, each relational table is assigned a fixed privacy budget with $\epsilon=1.5$ and $\delta=0.00005$ to release synopses. This setup ensures that the total privacy loss per table does not exceed $\epsilon=1.5$. For all equal-width histograms generated in SPECIAL, we configure them to have at most 8 bins. For Shrinkwrap, we use a privacy budget of $\epsilon=1.5$ and $\delta=0.00005$ for each query processing, as specified in (Bater et al., 2018). However, Shrinkwrap does not provide bounded privacy loss guarantees, so the privacy guarantee of Shrinkwrap will be no better than SPECIAL. We conduct all experiments on bare-metal Mac machines with M2 Max CPUs and 96GB unified memory. In addition, we consider Q2, Q4, Q8 are default testing queries in each query type, and are used as the representative workload to guide synopses. Moreover, for baseline systems, as they do not have join ordering optimizations, thus we will assume a random join order for them.
Figure 4. In-depth comparisons in execution plans: (i) The exhaustive padding in SMCQL can lead to significant memory blowup; (ii) Both SMCQL and Shrinkwrap suffer from suboptimal join ordering; (iii) Although Shrinkwrap reduces intermediate sizes, it still requires substantial memory to materialize the Cartesian-style standard oblivious join; (iv) SPECIAL effectively identifies optimal join orders and significantly reduces intermediate materialization sizes; (v) The IdxAcc method in SPECIAL not only bypasses secure selections but also directly reduces input I/O costs.
7.2. End-to-end performance comparisons
To address Question-1, we first conduct an end-to-end performance comparison of SPECIAL, Shrinkwrap, and SMCQL across the full benchmark workloads (Q1-Q8). We use two typical performance metrics, query execution time and memory usage in our evaluations. The results are summarized in Figure 5 and Figure 6. Due to significant memory consumption, SMCQL cannot complete the full benchmark and stops at Q4. Therefore, we only include full statistics up to Q3, and for Q4, we provide a projected performance evaluation using 10% of the input data.
Figure 5. End-to-end comparison: query latency
Figure 6. End-to-end comparison: memory usage
Observation 1. SPECIAL significantly outperforms all baseline systems in query latency, achieving performance gains of up to 3618.3×3618.3$\times$3618.3 × for linear queries, 114×114$\times$114 × for binary joins, and 80.3×80.3$\times$80.3 × for multi-joins. Figure 5 shows significant performance gains of SPECIAL in contrast to Shrinkwrap and SMCQL. For linear queries, SPECIAL exhibits substantial performance gains, achieving up to 3618.3×3618.3$\times$3618.3 × improvement (Q2), primarily due to its index-based fast data access. This approach not only reduces large I/O costs but also eliminates the need for secure computation by enabling direct data retrieval via DP indexes. We observe that in binary joins, SPECIAL has a less pronounced advantage over Shrinkwrap. This is because Shrinkwrap do not have sub-optimal planning issue when dealing with binary join—as there is only one join order. Hence, the performance gains for SPECIAL in Q3 and Q4 stem solely from fast data access and the efficient (MX)Join method. However, for more complex multi-way joins, SPECIAL’s advantage becomes more pronounced again due to its ability to select the optimal join order. For instance, Shrinkwrap shows more than 80×80$\times$80 × slow down in query latency than SPECIAL in group Q7.
Observation 2. SPECIAL shows profound improvement in memory usage (Figure 6) against baseline systems, especially in complex multi-way joins. This is primarily due to two factors: First, the (MX)JOIN used by SPECIAL is more memory-efficient compared to the joins implemented by Shrinkwrap and SMCQL. Second, SPECIAL’s capability to identify optimal execution plans significantly reduces total intermediate sizes, which is particularly beneficial for complex joins that suffer from sub-optimal or exhaustive padding in other systems. To better understand the substantial improvements SPECIAL achieves—for instance, up to 928.2×928.2$\times$928.2 × over Shrinkwrap and more than $10^{5}\times$ over SMCQL—we will zoom into a specific query, Q6, and compare the detailed execution plans of the three systems. The choice of Q6 is strategic because its complexity sufficiently highlights the differences in execution plans, yet it remains simple enough for clear visual representation. Our comparison features four execution plans: a plaintext optimal plan derived directly from the logical plan using actual cardinalities, illustrating the ground truth optimal execution; a hypothetical SMCQL plan (since Q6 cannot be completed by SMCQL so we project the cardinalities); and two actual execution plans from our experiments with Shrinkwrap and SPECIAL. The detailed comparisons and observations are summarized in Figure 4.
7.3. Privacy comparisons
Continue to address Question-1, we now compare the privacy guarantees between Shrinkwrap and SPECIAL, with the results detailed in Figure 7. As privacy loss is independent of query types, our analysis will focus on comparing the privacy composition curves under continual query answering for both Shrinkwrap and SPECIAL. We consider two privacy composition models: advanced composition (Adv.)(Dwork et al., 2014) and concentrated DP composition (CDP)(Bun and Steinke, 2016).
Figure 7. End-to-end comparison privacy loss.
Observation 3. Under continual query answering, SPECIAL demonstrates significantly lower privacy loss compared to Shrinkwrap, achieving up to 89.01×89.01$\times$89.01 × and 38.91×38.91$\times$38.91 × improvements in the Adv. and CDP modes, respectively. The privacy loss of SPECIAL is bounded to the initial synopsis release stage, so continual query answering does not incur additional privacy loss. In contrast, Shrinkwrap’s privacy loss accumulates over time as each new query allocates a fresh privacy budget. Consequently, its privacy loss exhibits a logarithmic growth, as shown in Figure 7. This accumulation can result in significant privacy degradation when processing a large number of queries. For example, answering 100 queries in Shrinkwrap could result in a privacy loss of $\epsilon>100$ in Adv. and $\epsilon\approx 60$ in CDP., respecitvely, even if each query only uses a small privacy budget of $\epsilon=1.5$. As such, SPECIAL demonstrate significant improve in privacy guarantees towards SOTA DPSCA.
7.4. Privacy efficiency tradeoffs
We address Question-2 by evaluating SPECIAL at various privacy levels and observing the performance impacts. Specifically, we maintain $\delta$ constant while varying $\epsilon$ from $0.1$ to $10$ and assess the performance across default testing queries (Q2, Q4, Q8). The results are shown in Figure 8 and Figure 9.
Figure 8. Privacy vs. Performance (query time)
Figure 9. Privacy vs. Performance (memory)
Observation 4. The privacy-efficiency tradeoff generally exists but exhibits varying trends at different privacy levels. For instance, SPECIAL shows a clear tradeoff at higher privacy levels ($\epsilon<1$), while at lower privacy levels ($\epsilon>1$), the tradeoff becomes less pronounced. When $\epsilon$ increases from $0.1$ to $1$, both memory usage and query latency for all test queries significantly decrease. However, increasing $\epsilon$ from $1$ to $10$ shows no significant performance gains. This may indicate that once $\epsilon$ exceeds $1$, the impact of noises on cardinality estimation or index building is alredy minimal, and further reductions in $\epsilon$ do not lead to notable improvements. Therefore, if high privacy protection is required, practitioners should carefully fine-tune privacy parameters to optimize performance. Conversely, if performance is the priority, setting $\epsilon$ near 1 is typically sufficient.
Observation 5. Different queries exhibit varying performance sensitivities to changes in privacy levels. Multi-join queries are significantly affected when privacy levels change. For instance, the performance overhead of Q8 at $\epsilon=0.1$ is more than 3×3$\times$3 × higher than at $\epsilon=1$. In contrast, linear query (Q2) and binary join (Q4) overhead increases to only about 1.5×1.5$\times$1.5 × under the same privacy settings. This difference stems from the error amplification inherent in multi-join operations, which frequently reuse synopses and require multiple cardinality estimations. Therefore, practitioners dealing with complex multi-join workloads may need careful fine-tune privacy parameters, as performance impacts can be substantial.
7.5. Scaling experiments
To address Question-3, we stress SPECIAL with scaled workloads by duplicating the raw financial dataset to sizes of 2×2$\times$2 ×, 4×4$\times$4 ×, and 8×8$\times$8 ×. We then evaluate the default testing queries on these scaled datasets to assess SPECIAL’s efficiency. The results are shown in Figure 10.
Figure 10. End-to-end comparison privacy loss.
Observation 6. SPECIAL demonstrates significant potential to scale up to multi-million records, even when handling complex 5-way joins. Figure 10 shows SPECIAL’s effective scaling: up to 8×8$\times$8 × data for linear queries and binary joins, and up to 4×4$\times$4 × data for complex 5-way joins like Q8. For instance, Q2 can be completed within 290ms under 8×8$\times$8 ×, and in fact, since selection is bypassed (due to index access), thus the cost is mainly on I/O costs. Q4 finishes in 289 minutes at the same scale 8×8$\times$8 ×, while the more complex 5-way join Q8 takes less than 280 minutes for 4×4$\times$4 × data. As a reference, Shrinkwrap would require over 1035 minutes to complete Q8 even with unscaled data (Figure 5). This result can, from another angle, showcase the significant performance improvements of SPECIAL over current SOTA DPSCA designs.
8. Discussion
In this section, we discuss potential extensions to accommodate broader data models and possible adaptations of SPECIAL designs to other secure primitives.
Supporting growing data. In general SPECIAL considers read-intensive analytical workloads and and a static data model, which is the same as current SOTA SCA efforts (Poddar et al., 2021; Bater et al., 2018; Liagouris et al., 2023; Volgushev et al., 2019). In this section, we will briefly discuss how SPECIAL could be extended to support the growing data model where new records are continuously added to the logical database (Wang et al., 2022, 2021; Zhang et al., 2023). One possible strategy involves segmenting the large logical database into discrete, non-overlapping time intervals, each represented by its own database with independent synopses. Specifically, for new incoming data, SPECIAL will initially cache it until a designated time interval concludes. After this period, SPECIAL will batch all the cached data into a new database corresponding to the past time interval and release synopses for this database. As the data within these intervals is disjoint, queries can be easily partitioned and independently optimized and executed across corresponding databases in parallel, before being aggregated. Note that prior to releasing new synopses, all cached data will be processed using standard SCA processing (e.g. SMCQL). Since each synopses release covers only newly inserted data, the total privacy loss adheres to parallel composition (Dwork et al., 2014), and thus the bounded privacy loss still hold.
Adaptability to other secure primitives. Beyond MPC-based SCA, the key design insights from SPECIAL are transferable to other secure outsourced databases such as the TEE-based analytical systems (Qiu et al., [n.d.]; Qin et al., 2022; Eskandarian and Zaharia, 2017; Zheng et al., 2017) or searchable encryption based private key-value store (Amjad et al., 2021; Groce et al., 2019; Chen et al., 2018). This is because SPECIAL’s core components, including synopses management, indexing, query planning, and compaction estimation, can all be executed without relying on secure computation primitives. For example, in TEE databases, one may conduct a similar query planning outside the TEE, with the optimal plan then executed within the TEE for fast oblivious processing. Similarly, SPECIAL’s design principles could be applied to searchable encryption to construct private indexes in advance on sorted data, enable fast and private information retrieval.
9. Related Work
SCA systems. Two main approaches exist for designing MPC-based SCA systems. The first is peer-to-peer paradigm (Aggarwal et al., 2005; Mouris et al., 2024; Poddar et al., 2021; Volgushev et al., 2019; Wang and Yi, 2021), where the goal is to improve efficiency by decomposing analytical queries and pushing them to data owners, so that they can either directly process in clear or running MPCs across a smaller group of parties. Unfortunately, this approach burdens data owners with heavy computational overheads, particularly for complex operations like joins. Moreover, given real-world data owners often lack robust computing resources and reliable service capabilities, these methods are difficult to scale and promise consistent SCA services to external analysts. On the other hand, the server-aided-MPC model (Wang et al., 2021; Bater et al., 2018, 2020; Wang et al., 2022; Zhang et al., 2023; Roy Chowdhury et al., 2020; Kamara et al., 2011; Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023; Mohassel and Zhang, 2017) offers a more practical solution. It allows data owners to outsource both expensive MPC computations and secure data storage to a set of capable cloud service providers (CSPs), who then jointly evaluate MPC to provide reliable SCA services. For the SPECIAL prototype, we focus on the server-aided-MPC model under a strong corruption assumption, allowing up to $n-1$ of $n$ servers to be compromised and collate with others. While, as we discussed before, SPECIAL’s core design principles are protocol-agnostic. This enables interoperability with various MPC models, including the peer-to-peer settings or a weaker corruption setting (for performance purpose) where a supermajority of servers need to be honest (Tan et al., 2021; Knott et al., 2021; Liagouris et al., 2023).
DP leakages. Leakage-abuse attacks (Cash et al., 2015; Kellaris et al., 2016; Blackstone et al., 2019; Oya and Kerschbaum, 2021; Grubbs et al., 2018; Shang et al., 2021; Zhang et al., 2016), exploit data-dependent processing patterns, are persistent threats to SCA systems. To mitigate these risks, oblivious computation techniques (Keller and Scholl, 2014; Pinkas and Reinman, 2010; Stefanov et al., 2011; Chang et al., 2022; Mishra et al., 2018; Crooks et al., 2018; Jönsson et al., 2011; Sasy et al., 2022; Asharov et al., 2020; Wang et al., 2014; Pettai and Laud, 2015; Wang et al., 2023; Tinoco et al., 2023) have become the de facto solution—processing data such that the entire execution transcript (e.g. memory traces and read/write volumes) is padded to a worst-case constant. While these techniques ensure the strongest privacy guarantees by eliminating data-dependent leakages, they also introduce a fundamental contention with modern database optimizations, which often rely heavily on data-dependent operations (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022). To this end, many recent efforts seek a practical balance in the privacy-performance trade-off by allowing controlled leakage under DP (Bater et al., 2018; Wang et al., 2021; Wagh et al., 2021; Wang et al., 2022; Groce et al., 2019; He et al., 2017; Zhang et al., 2023; Qin et al., 2022; Chan et al., 2022; Wagh et al., 2016; Chu et al., 2021; Qiu et al., [n.d.]; Patel et al., 2019). However, a common issue of these approaches is unbounded privacy loss. While some works propose to address this (Zhang et al., 2023; Bogatov et al., 2021; Qiu et al., [n.d.]), their approaches are restricted to only simple linear queries. Furthermore, existing DP leakage designs also lack lossless guarantees for complex query processing. The inherent randomness of DP noises (Wang et al., 2021; Groce et al., 2019; Wang et al., 2022) and uncertainty in complex join sensitivities (Chu et al., 2021; Bater et al., 2018; Wang et al., 2022) often lead to improper output compaction and lose tuples. SPECIAL addresses all these limitations together, and to our knowledge, is the first system to ensure both bounded privacy and lossless results for securely processing complex queries.
SCA query planning. Query planning (Silberschatz et al., 2011) is crucial in conventional databases. A few studies (Bater et al., 2017; Volgushev et al., 2019; Poddar et al., 2021; Liagouris et al., 2023) explored query planning within SCA frameworks. Unlike conventional planners, which exploit size disparities across different execution plans to choose the most efficient one with minimized plan sizes (Elmasri et al., 2020; Harmouch and Naumann, 2017; Han et al., 2021), these methods typically prohibit the use of such data-dependent information. Consequently, they primarily rely on data-independent metrics for planning. For example, (Bater et al., 2017; Liagouris et al., 2023; Volgushev et al., 2019) optimize based on the MPC execution costs, while (Poddar et al., 2021) focuses on decomposing large MPC tasks into smaller 2/3PC ones. Since these methods do not optimize or compact intermediate sizes, significant query execution overhead from these large sizes remains evident, especially when processing complex multi-joins (Bater et al., 2018; Liagouris et al., 2023). Shrinkwrap (Bater et al., 2018) uses DP leakage models and allows planner to optimize query intermediate sizes in a privacy-preserving manner. However, it cannot estimate cardinalities before runtime, mandating a random selection of plan structures (e.g. join order) before execution. Consequently, planning focuses solely on optimally allocating privacy budgets across the chosen plan, which often result in suboptimal query executions. In contrast, SPECIAL offers an advanced query planner capable of pre-estimating plan sizes and comparing costs among different plan structures before execution.
10. Conclusion
We introduce SPECIAL, the first SCA system that simultaneously supports: (i) handling complex queries with bounded privacy loss; (ii) advanced query planning that effectively exploit plan intermediate sizes before runtime; and (iii) delivering exact query results without missing tuples. This is achieved through a novel synopses-assisted SCA design, where a set of private table statistics are released with one-time privacy cost to guide subsequent secure SCA planning and processing. The benchmark results demonstrate that SPECIAL significantly outperforms SOTA SCA systems and scales complex multi-join queries to multi-million datasets
Appendix A Privacy Proof
Theorem A.1.
The one-sided Laplace mechanism as described in Definition 4.1 satisfies $(\epsilon,\delta)$-DP.
Proof.
WLOG. we illustrate proof details by assuming positive mechanism, $\mathsf{Lap}^{+}$, and by symmetric, the properties proved here also holds for $\mathsf{Lap}^{-}$. In general, we write the mechanism as $\mathcal{M}(D)\leftarrow f(D)+\mathsf{Lap}^{+}(\epsilon,\delta)$, where $f$ is a counting query. Note that the sensitive for counting query is 1, and thus we will implicitly omit the sensitivity notations.
Given any two neighboring databases $D$ and $D^{\prime}$, and let $c\leftarrow f(D)$ and $c^{\prime}\leftarrow f(D^{\prime})$, WLOG. we will assume $c\leq c^{\prime}$. Now we will evaluate three disjoint output domains for $f$, namely $O_{1}=(-\infty,c)$, $O_{2}=[c,c^{\prime}]$, and $O_{3}=(c^{\prime},+\infty)$. Since the noise is non-negative, thus it’s clear that $\textup{Pr}[\mathcal{M}(D)\in O_{1}]=0$ and $\textup{Pr}[\mathcal{M}(D^{\prime})\in O_{1}]=0$. Next, we evaluate the domain $O_{3}$ and consider an auxiliary domain $O^{*}=\{x-c^{\prime}:x\in O_{3}\}$. We compute the following ratio:
(6) $\begin{split}\frac{\textup{Pr}[\mathcal{M}(D)\in O_{3}]}{\textup{Pr}[\mathcal{M}(D^{\prime})\in O_{3}]}&=\frac{\textup{Pr}[c+\mathsf{Lap}^{+}(\epsilon,\delta)\in O_{3}]}{\textup{Pr}[c^{\prime}+\mathsf{Lap}^{+}(\epsilon,\delta)\in O_{3}]}\\ &=\frac{\textup{Pr}[\mathsf{Lap}^{+}(\epsilon,\delta)\in O^{*}]}{\textup{Pr}[\mathsf{Lap}^{+}(\epsilon,\delta)\in O^{*}+(c^{\prime}-c)]}\\ &=\frac{e^{-\epsilon|x-\mu|}}{e^{-\epsilon|x-(c^{\prime}-c)-\mu|}}\\ &=e^{\epsilon(|x-(c^{\prime}-c)-\mu|-|x-\mu|)}\\ &\leq e^{\epsilon(|c^{\prime}-c)|)}~{}~{}(\textup{triangle~{}inequality})\\ &=e^{\epsilon}\end{split}$
Then we evaluate the last output domain $O_{2}$, and we can see that $\textup{Pr}[\mathcal{M}(D^{\prime})\in O_{2}]=0$ and $\textup{Pr}[\mathcal{M}(D)\in O_{2}]\geq 0$, and thus there will be unbounded privacy loss within the domain of $O_{2}=[c^{\prime},c]$. However, we can bound this probability by $\delta$, which is equivalent to prove that with density function
$\textup{Pr}[z=x]=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon|x-\mu|},$
where $\mu=1-\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))$, the probability of sampling a value $z$ within the range $(-\infty,1]$ is less than or equal to $\delta$. Since we consider a discrete distribution, thus we compute the following
(7) $\begin{split}\Pr[z<1]&=\sum_{x=-\infty}^{0}\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon(\mu-x)}\\ &=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon\mu}\sum_{x=-\infty}^{0}e^{\epsilon x}\\ &=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon\mu}\frac{1}{1-e^{-\epsilon}}\quad\text{(geometric series sum of }e^{\epsilon x})\\ &=\frac{e^{\epsilon}-1}{e^{\epsilon}+1}e^{-\epsilon\left(1-\frac{1}{\epsilon}\ln(\delta(e^{\epsilon}+1))\right)}\frac{1}{1-e^{-\epsilon}}\quad\text{(substitute }\mu)\\ &=\frac{e^{\epsilon}}{e^{\epsilon}+1}e^{-\epsilon+\ln(\delta(e^{\epsilon}+1))}\\ &=\frac{e^{\epsilon}}{e^{\epsilon}+1}e^{-\epsilon}\delta(e^{\epsilon}+1)\\ &=\delta.\end{split}$
Combining all these we obtain the following
(8) $\begin{split}\textup{Pr}[\mathcal{M}(D)\in O]&=\textup{Pr}[\mathcal{M}(D)\in(O_{1}\cup O_{2}\cup O_{3})]\\ &=\textup{Pr}[\mathcal{M}(D)\in O_{2}]+\textup{Pr}[\mathcal{M}(D)\in O_{3}]\\ &\leq\delta+e^{\epsilon}\textup{Pr}[\mathcal{M}(D^{\prime})\in(O_{2})]\end{split}$
The claim thus holds. ∎
Next we prove Theorem 4.3.
Proof.
(Theorem 4.3) We prove this theorem by analyzing privacy loss using composition theorems. Generating noisy histograms with one-sided noise is equivalent to releasing multiple counting queries across disjoint datasets. Therefore, by the parallel composition theorem (Dwork et al., 2014), each histogram release is $(\epsilon,\delta)$-DP. The release of the maximum frequency via a noisy max mechanism is $(\epsilon,0)$-DP. Given $c$ pairs, the sequential composition theorem (Dwork et al., 2014) applies, resulting in a total privacy loss under $2c$-fold advanced composition. Consequently, the entire process adheres to $(\hat{\epsilon},\hat{\delta})$-DP, with $\hat{\epsilon}\leq 6\epsilon\sqrt{c\ln(1/\delta)}$.
Theorem A.2.
The privacy of SPECIAL adheres to Definition 3.1
Proof.
Since the entire DP synopses releasing satisfies DP. Thus, we prove this theorem by constructing a simulator, $\mathcal{S}$, which takes only the DP synopses as input and can simulate execution transcripts that are indistinguishable from those of the actual protocol. For ease of notation, we abstract SPECIAL as a composed secure protocol $\pi$.
We first examine a hybrid protocol, $\pi_{1}$, differing from $\pi$ in that the secure outsourcing procedure—where owners upload secret shares of their data to two servers—is replaced by a simulator, $\mathcal{S}_{1}$. This simulator, which takes no input, simulates secret shares with randomly sampled data $(x_{1},x_{2})\xleftarrow[]{\mathsf{rd}}\mathbb{Z}_{2^{32}}$. We say that a $\mathcal{S}_{1}$ must exist that produces results indistinguishable from the secret shares in $\pi$, as the absence of such a simulator would imply the non-existence of secure secret sharing techniques. Next, we consider another hybrid protocol, $\pi_{2}$, which differs from $\pi_{1}$ only in that the query planning function is replaced by a simulator $\mathcal{S}_{2}$. This simulator takes as inputs all synopses and a set of random values (simulated secure outsourced data) generated by $\mathcal{S}_{1}$, and produces a simulated execution plan for every requested query. We say that an $\mathcal{S}_{2}$ should also exist with outputs indistinguishable from $\pi$’s execution plans. Given that Selinger optimizers can pre-generate execution plans using only table statistics, the non-existence of such a simulator, $\mathcal{S}_{2}$, would imply the non-existence of Selinger optimizers, which is contradictory. Finally, we construct $\mathcal{S}=(\mathcal{S}_{1},\mathcal{S}_{2})$, where $\mathcal{S}_{1}$ simulates secure data outsourcing and $\mathcal{S}_{2}$ simulates query planning. $\mathcal{S}$ executes the query plan using randomly generated values from $\mathcal{S}_{1}$ through MPC computation. Since both the execution plans and input data are indistinguishable from those in $\pi$, the MPC execution must also be indistinguishable. Otherwise, it would violate the security guarantees of MPC. ∎
Appendix B Query Workloads
We illustrate the benchmark query workloads as follows:
Listing 1: Benchmark query workload